User Tools

Site Tools


lbaops:lbamar2008:v252dcdlog

Setup:

RCP into Chan 2 on Frequency Translator, LCP into Chan 1

Sky Freq 22316, Agilent 17800, Fixed 4100 = IF 416, SMY LO 736

DAS profile = VSOP_HO.PRO

used 25th harmonics for coherence at 892.2 (→22305) and 892.84 (→22321)

Recorded to disks SWIN V019A, labels SWIN_V019A_P1 and SWIN_V019A_P2

We had a few wind stows early on in the experiment, but was fine from then on.

Problems: After the experiment we saw that the LCP was subject to many 'dropouts'. We later tracked this down to a faulty cable. The weather was terrible for 22 GHz at the start so we were unable to see the problem until the end when the weather had cleared and we could do some calibration scans. We suspect that the LHC would have been subject to 'dropouts' throughout the exp.

lbaops/lbamar2008/v252dcdlog.txt · Last modified: 2015/12/18 16:38 (external edit)